Session: Jun 2014 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Zelenchukskaya (Ptitcin A. at Wed Jun 04 00:36:29 2014) Success:
Scans from No0030 to No0044 missed due to participation in intensive geo session

Svetloe (Andreeva T. at Wed Jun 04 00:37:42 2014) Minor failures: WIND
Scans No0015- No0026 failed due to thunderstorm.

Badary (Zhukova at Wed Jun 04 05:37:39 2014) Success:
Pause for intensive geo session (scans no0038-1942)

Noto (Salvo Buttaccio at Wed Jun 04 05:46:32 2014) Minor failures: RECO
mk5 bank switch failure. About 1 hour lost.

Hartebeesthoek (Jonathan Quick at Wed Jun 04 07:00:52 2014) Success:
No known problems apart from the usual RFI in the lowest two sub-bands

Torun (K. Borkowski at Wed Jun 04 08:58:40 2014) Success:
No known problems.

Westerbork (Geert Kuper at Wed Jun 04 10:05:15 2014) Minor failures: PHAS
From 14:00 - 16:53 UT RT9 gave no signal but was not excluded from the adding box. After 16:53 excluded.

Jodrell_Bank (Jimmy Cullen at Wed Jun 04 11:59:10 2014) Success:
Due to a bogie failure, the Lovell was replaced by the MkII for this observation. Otherwise no problems.

Onsala (Jun Yang at Thu Jun 05 06:06:12 2014) Success:
No known problems.

Effelsberg (Uwe Bach at Thu Jun 05 13:15:42 2014) Minor failures: LATE
Start 1h late, because of focus control problems.

Robledo (Cristina Garcia Miro at Tue Jul 15 12:40:55 2014) Minor failures: TSYS LATE
DVP configuration script not properly defined regarding channels definition (channels 5 and 7) and IF input. Problem corrected in real-time. As a result 2 first sources not properly recorded (scans 64&65). All channels (1-8) injected to L-LCP (only pol available). L-band noise diode red, no tsys data available. Mark5C VDIF recording, label=JPL-0050/4000